OpenStudy (anonymous):

line segment RS is perpendicular to line segment PQ, and the coordinates are R(4,-5) S(-8,4) P(0,6) and Q(-3,y) what is the value of y ?

OpenStudy (anonymous):

Use slope1 * slope2 = -1 .... for 2 perpendiucular lines here slope of RS = (-5 - 4)/(4-(-8)) = - 9/12 = - 3/4 slope of PQ = (6 - y)/(0+3) = (6-y)/3 multiply the two to get (y-6)/4 = -1 y =2

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

thanks for giving me a thanks :P

OpenStudy (anonymous):

Lol